Sol Summary – September 28th

Sol: 1

Summary Title: The Orange Bucket

Author’s name: Jin Sia (HSO)

Mission Status: Nominal, Hab fully prepared for in-sim operations

Sol Activity Summary: We started the day with breakfast pancakes made
by Dave, during which we discussed how our daily routine will run for
the next few sols. Then, we got to work cleaning the Hab lower deck
and the RAM. After a thorough vacuuming and mopping, including Inga
and Lindsay going over every square inch of the lower deck with paper
towels, the floors are now sparkling clean. We hope they won’t get too
dirty again if we keep our EVA boots confined to the airlocks –
perhaps this is a bit optimistic, given that we are surrounded by
literal mountains of dust! After a lunch of broccoli cheddar soup made
by Lindsay, we tested out the EVA suits by donning and doffing them on
the lower deck, then walking through the tunnel between the Hab and
the RAM. We found that the old style of spacesuits worked best for us
if the straps are secured properly. However, a ferocious Martian dust
storm (that contained a suspiciously large amount of liquid water)
quickly forced us back into the Hab. I proceeded to make a Malaysian
staple, Chinese-style fried rice. It turned out to be the best fried
rice I’ve ever made – does this mean that restaurants use freeze-dried
ingredients for fried rice?! Overall, it was a productive day that
ended with us all being very well-fed. We may find we’ve gained weight
at the end of this…

Look Ahead Plan: Inga will begin her psychological journal pilot study
later tonight. We also hope to file a plan for our first two EVAs on
the Red Planet tomorrow, if weather permits.

Anomalies in work: We encountered issues with the new spacesuit neck
rings that made them uncomfortable to wear. We will contact Dr. Rupert
to discuss next steps.

Weather: Warm and sunny this morning, but a thunderstorm came near the
Hab late in the afternoon. We did not need to evacuate to the lower

Crew Physical Status: All crew nominal, feeling better than yesterday.

EVA: None today.

Reports to be filed: GreenHab, Science, Photo, EVA request

Support Requested: RAM seems to be missing a smoke detector,
requesting an additional one to put there.


Copyright © The Mars Society. All rights reserved. | Main Site